OpenStudy (anonymous):

Two sides of a triangle have lengths of 12m and 15m the angles between them are increasing at a rate of 2 degrees per minute.How fast is the length of the third side increasing when the angle between the sides of fixed length is 60 degrees? Please help I am confused.

OpenStudy (anonymous):

@ajprincess @hartnn please help

hartnn (hartnn):

u need cosine rule here, know what is it ?

OpenStudy (anonymous):

no

hartnn (hartnn):

 from cosine-rule, we have \(\huge c^2=a^2+b^2-2ab\cos\theta\)

OpenStudy (anonymous):

okay

hartnn (hartnn):

Given , a,b,\(\frac{d}{dt}\theta\)

hartnn (hartnn):

and \(\theta\)

hartnn (hartnn):

u could identify all 4 of these values, right ?

OpenStudy (anonymous):

im not sure

hartnn (hartnn):

two sides, a and b are given as ....? the rate of angle change per minute = \(\frac{d}{dt}\theta\)= ....? and the angle itself is given as \(\theta=....degrees?\)

OpenStudy (anonymous):

okay

hartnn (hartnn):

so tell me all 4 values, first.

hartnn (hartnn):

next i need to know whether u can differentiate?

OpenStudy (anonymous):

im confused

hartnn (hartnn):

on which value ?

OpenStudy (anonymous):

you showed me only 3

hartnn (hartnn):

2 sides, a and b are 2 values.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

never mind i already got the answer thanks anyway